# Current affiliation: ]Fakultät für Physik, Technische Universität Dortmund, 44227 Dortmund, Germany Spin Liquid Mimicry in the Hydroxide Double Perovskite CuSn(OD)6 Induced by Correlated Proton Disorder

### v1

@@ III Results and Discussion
The magnetic entropy obtained by integrating is shown in Fig. 2 (c). A point at (0,0) was added to all curves for an approximate extrapolation to ; this extrapolated region accounts for 0.13 J mol-1 K-1. The magnetic entropy saturates around 30 K, far above the specific-heat anomalies as expected for frustrated spin systems, but it asymptotes only 85% of . This suggests that not all spin degrees of freedom are available at low temperatures. This may indicate a magnetic transition at temperatures lower than our minimum of 0.45 K, a model subtraction issue, or quantum entanglement suppressing classical thermal excitations.
@@
Taking all our data together, we see no evidence of even short-range magnetic order down to 50 mK, despite a Curie-Weiss temperature of 7.1 K and an apparent interaction energy scale of 18 T. The nearly 2.5 orders of magnitude separating the temperature of our neutron measurement from the interaction energy scale suggests very strong frustration in CuSn(OD)6. However, the suppression of magnetic order may also arise from quantum fluctuations. The Mn compound has a similar but higher-symmetric crystal structure with analogous proton disorder, but its magnetic moments are classical spins which are not expected to experience significant quantum fluctuations. The lower symmetry in the Cu analog would ordinarily be expected to partially relieve frustration, making long-range order more stable, so the lack of order suggests that quantum fluctuations play a key role in CuSn(OD)6.
@@
In both herbertsmithite and YbMgGaO4, spin-liquid mimicry arises from quenched substitutional disorder that is difficult to control. In CuSn(OD)6, however, the disorder arises from ice-rule physics among the hydrogen atoms which fill the perovskite A site. In water ice this can be tuned with pressure, leading to a rich phase diagram, and our previous neutron diffraction pressure study indicated that the hydrogen sites in CuSn(OD)6 are likely to merge at pressures on the order of 20 GPa [38]. This is an experimentally-accessible pressure. It should be possible to continuously tune the correlated hydrogen disorder out of existence while performing a variety of measurements, replacing the spin-liquid-like ground state we observe with either magnetic order or a quantum spin liquid.
@@ IV Conclusions
Samples and data are available upon reasonable request from D. C. Peets or D. S. Inosov; data underpinning this work is available from Refs. 47, 48.
@@ Acknowledgements.
- Parui et al. [2025] K. K. Parui, A. A. Kulbakov, E. Häußler, N. S. Pavlovskii, A. Mannathanath Chakkingal, M. Avdeev, R. Gumeniuk, S. Granovsky, A. Mistonov, S. A. Zvyagin, T. Doert, D. S. Inosov, and D. C. Peets, Disordered ground state in the 3D face-centred frustrated spin- system MnSn(OH) (2025), arXiv:2502.12433 [cond-mat.str-el] .

### v2

@@ III Results and Discussion
The magnetic entropy obtained by integrating is shown in Fig. 2 (c). A point at (0,0) was added to all curves for an approximate extrapolation to ; this extrapolated region accounts for 0.13 J mol-1 K-1. The magnetic entropy saturates around 30 K, far above the specific-heat anomalies as expected for frustrated spin systems, but it asymptotes only 85% of . This suggests that not all spin degrees of freedom are available at low temperatures. This may indicate residual quantum fluctuations, a magnetic transition at temperatures lower than our minimum of 0.45 K, a model subtraction issue, or quantum entanglement suppressing classical thermal excitations.
@@
Taking all our data together, we see no evidence of even short-range magnetic order down to 50 mK, despite a Curie-Weiss temperature of 7.1 K and an apparent interaction energy scale of 18 T. The nearly 2.5 orders of magnitude separating the temperature of our neutron measurement from the interaction energy scale suggests very strong frustration in CuSn(OD)6. However, the suppression of magnetic order may also arise from the strong quantum fluctuations expected in Cu materials. The Mn compound has a similar but higher-symmetric crystal structure with analogous proton disorder, but its magnetic moments are classical spins which are not expected to experience significant quantum fluctuations. The lower symmetry in the Cu analog would ordinarily be expected to partially relieve frustration, making long-range order more stable, so the lack of order suggests that quantum fluctuations play a key role in CuSn(OD)6.
@@
In both herbertsmithite and YbMgGaO4, spin-liquid mimicry arises from quenched substitutional disorder that is difficult to control. In CuSn(OD)6, however, the disorder arises from ice-rule physics among the hydrogen atoms which fill the perovskite A site. Figures 4(a) and 4(b) show two possible hydrogen configurations in a selected A site. These should be degenerate and completely decoupled from the other A sites. An example nearest-neighbor exchange pathway is shown in Fig. 4(c), where all hydrogen positions are partially occupied. Routes that do not pass through disordered hydrogen bonds traverse oxygen atoms to which disordered hydrogen atoms are bonded. In water ice, hydrogen disorder can be tuned with pressure [47], leading to a rich phase diagram, and our previous neutron diffraction pressure study predicted that the hydrogen sites in CuSn(OD)6 would merge at pressures on the order of 20 GPa [38], at which point the hydrogen disorder would be quenched. This is an experimentally-accessible pressure. It is unlikely that all hydrogen sites would merge at the same pressure, potentially leading to a cascade of structural phases as seen in water ice. It should be possible to tune the correlated hydrogen disorder out of existence while performing a variety of measurements, replacing the spin-liquid-like ground state we observe with either magnetic order or a quantum spin liquid.
@@ IV Conclusions
Samples and data are available upon reasonable request from D. C. Peets or D. S. Inosov; data underpinning this work is available from Refs. 48, 49.
@@ Acknowledgements.
- Parui et al. [2025] K. K. Parui, A. A. Kulbakov, E. Häußler, N. S. Pavlovskii, A. Mannathanath Chakkingal, M. Avdeev, R. Gumeniuk, S. Granovsky, A. Mistonov, S. A. Zvyagin, T. Doert, D. S. Inosov, and D. C. Peets, Disordered ground state in the 3D face-centred frustrated spin- system MnSn(OH), Phys. Rev. B 112, (2025), arXiv:2502.12433 [cond-mat.str-el] .
- Komatsu [2022] K. Komatsu, Neutrons meet ice polymorphs, Crystallogr. Rev. 28, 224 (2022).